If a uniform bar is to balance on a fulcrum, the ration of the larger weight to the smaller weight must equal the ratio of the smaller distance to the larger distance.
If an 8 kg weight is 7.2m from the fulcrum, how far from the fulcrum is a 9 kg weight which balances it?
HOLY TOLEDO THIS QUESTION IS RANK. PLEASE HELP

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(12126)



Just a simple proportion will do the trick:



Cross multiply:



You can do your own arithmetic.


John


Answer by stanbon(48535)

If a uniform bar is to balance on a fulcrum, the ratio of the larger weight to the smaller weight must equal the ratio of the smaller distance to the larger distance.
If an 8 kg weight is 7.2m from the fulcrum, how far from the fulcrum is a 9 kg weight which balances it?
---------------------------------------
x/9 = 7.2/8
x = 9(7.2/8)
x = 9*0.9
x = 8.1 meter
======================
Cheers,
Stan H.